# A unified convention for achievement positional games

### v1

@@ 4.4 AchievementPos(3,3) is PSPACE-complete
Input : A logic formula $\phi$ in CNF form, with clauses of size exactly 3, and an even number of variables with a prescribed ordering $x_{1},\ldots,x_{2n}$.
Output : T if Satisfier has a winning strategy, F otherwise.
Stockmeyer and Meyer [SM73] have shown that 3-QBF is PSPACE-complete. We now prove the PSPACE-hardness of AchievementPos($3$,$3$) using a reduction from 3-QBF.
Theorem 4.6.
AchievementPos($3$,$3$) is PSPACE-complete.
Proof.
Let $\phi$ be a logic formula in CNF form, with clauses of size exactly 3 and variables $x_{1},\ldots,x_{2n}$. We are going to build an achievement positional game $\mathcal{G}$ with edges of size at most 3 such that Left has a winning strategy on $\mathcal{G}$ as second player if and only if Falsifier has a winning strategy on $\phi$. Before giving the formal definition of $\mathcal{G}$, let us explain the idea. We see Left as Falsifier and Right as Satisfier. For each variable $x_{i}$, we will create (among others) four vertices $t_{i,R}$, $f_{i,R}$, $t_{i,L}$, $f_{i,L}$. There will be two phases in $\mathcal{G}$.
  • Phase 1: Both players take turns choosing valuations for the variables $x_{i}$, in increasing order of $i$. More precisely:
    • For all odd $i$, Right will be forced at some point to pick either $t_{i,R}$ or $f_{i,R}$, which corresponds to her setting $x_{i}$ to T or F respectively.
    • For all even $i$, Left will be forced at some point to pick either $t_{i,L}$ or $f_{i,L}$, which corresponds to him setting $x_{i}$ to T or F respectively.
    Additionally to the edges that will be there to force both players’ moves, there will be blue clause-edges corresponding to the clauses of $\phi$. The important thing is that the clause-edges will only contain $t_{i,R}$’s and $f_{i,R}$’s, so that Left never plays inside them during Phase 1: this ensures that he does not create threats with the clause-edges, so that his moves are indeed forced. If the valuation $\mu$ built by both players satisfies $\phi$, then Right will have picked at least one vertex in every clause-edge by the end of Phase 1, thus ensuring that she will draw the game.
  • Phase 2: Left picks the complementary vertices to those picked by Right during Phase 1. This means that, for all $i$, Left picks $t_{i,R}$ (resp. $f_{i,R}$) if Right had picked $f_{i,R}$ (resp. $t_{i,R}$). This way, if $\mu$ does not satisfy $\phi$, then Left will fill a clause-edge during Phase 2, thus winning the game.
Let us now define $\mathcal{G}=(V,E_{L},E_{R})$ rigorously.
  • We define $V=\bigcup_{1\leq i\leq 2n}(V_{i}\cup\{a_{i},b_{i},b^{\prime}_{i},c_{i},c^{\prime}_{i}\})\cup\{w\}$ where:
    • for all $1\leq i\leq 2n$: $V_{i}=\{t_{i,R},f_{i,R},t_{i,L},f_{i,L},u_{i},v_{i}\}$;
    • for all $1\leq i\leq 2n$: the vertices $a_{i}$, $b_{i}$, $b^{\prime}_{i}$, $c_{i}$ and $c^{\prime}_{i}$ are only relevant during Phase 2;
    • $w$ is an additional vertex.
  • We define $E_{R}=\bigcup_{1\leq i\leq 2n}E_{R}^{(i)}$ and $E_{L}=\bigcup_{1\leq i\leq 2n}E_{L}^{(i)}\cup\{e_{w}\}\cup\widetilde{E_{L}}\cup E_{L}^{\,\phi}$ where:
    • The edges in $E_{R}^{(i)},E_{L}^{(i)}$ are here to force both player’s moves during Phase 1. For convenience, we introduce the following notation. When the notation $e^{*}$ appears inside the definition of $E_{R}^{(i)}$, it means we add the edge $e$ if $i=1$, or the two edges $e\cup\{t_{i-1,R}\}$ and $e\cup\{f_{i-1,R}\}$ if $i\geq 2$. The point is that Right will pick one of $t_{i-1,R}$ or $f_{i-1,R}$, at which point the updated red edge will be $e$. Similarly, when the notation $e^{*}$ appears inside the definition of $E_{L}^{(i)}$, it means we add the edge $e$ if $i=1$, or the two edges $e\cup\{t_{i-1,L}\}$ and $e\cup\{f_{i-1,L}\}$ if $i\geq 2$.
      For odd $i$, we define $E_{R}^{(i)}=\{\{t_{i,R},f_{i,L}\}^{*},\{f_{i,R},t_{i,L}\}^{*},\{f_{i,R},f_{i,L},u_{i}\},\{t_{i,R},t_{i,L},v_{i}\}\}$ and $E_{L}^{(i)}=\{\{t_{i,R},f_{i,R}\}^{*},\{t_{i,R},u_{i}\}^{*},\{f_{i,R},v_{i}\}^{*},\{t_{i,L},u_{i}\}^{*},\{f_{i,L},v_{i}\}^{*}\}$.
      For even $i$, we define $E_{L}^{(i)}=\{\{t_{i,R},f_{i,L}\}^{*},\{f_{i,R},t_{i,L}\}^{*},\{t_{i,L},f_{i,L},u_{i}\},\{t_{i,L},f_{i,L},v_{i}\}\}$ and $E_{R}^{(i)}=\{\{t_{i,L},f_{i,L}\}^{*},\{t_{i,L},u_{i}\}^{*},\{f_{i,L},v_{i}\}^{*},\{t_{i,R},u_{i}\}^{*},\{f_{i,R},v_{i}\}^{*},\{t_{i,R},t_{i,L},v_{i}\},\{f_{i,R},f_{i,L},u_{i}\}\}$.
    • The blue edge $e_{w}=\{u_{2n},v_{2n},w\}$ is here to force Right to waste a move as we transition from Phase 1 to Phase 2.
    • The blue edges in $\widetilde{E_{L}}=\bigcup_{1\leq i\leq 2n}\{\{t_{i,R},a_{i},b_{i}\},\{t_{i,R},a_{i},c_{i}\},\{f_{i,R},a_{i},b^{\prime}_{i}\},\{f_{i,R},a_{i},c^{\prime}_{i}\}\}$ form butterflies which are here to force Right’s moves during Phase 2.
    • The blue clause-edges in $E_{L}^{\,\phi}$ are defined as follows. To every literal $\ell_{i}\in\{x_{i},\overline{x_{i}}\}$, we associate a vertex $v(\ell_{i})\in V$: if $i$ is odd then $v(x_{i})=t_{i,R}$ and $v(\overline{x_{i}})=f_{i,R}$; if $i$ is even then $v(x_{i})=f_{i,R}$ and $v(\overline{x_{i}})=t_{i,R}$. For each clause $\ell_{i_{1}}\vee\ell_{i_{2}}\vee\ell_{i_{3}}$ in $\phi$, we add the clause-edge $\{v(\ell_{i_{1}}),v(\ell_{i_{2}}),v(\ell_{i_{3}})\}$.
For visual aid, the game obtained for $n=2$, without the clause edges, is shown in Figure 10. Now, let us study how the game unfolds.
  1. 1)
    An important remark is that $\{t_{1,R},f_{1,L}\}$ and $\{f_{1,R},t_{1,L}\}$ are the only red edges of size 2 at the beginning of the game. We claim that Right is forced to pick $t_{1,R}$ or $f_{1,R}$ as her first move. Indeed, if she picks $t_{1,L}$, then Left picks $f_{1,R}$, and Right cannot defend the double threat of the blue edges $\{t_{1,R},f_{1,R}\}$ and $\{f_{1,R},v_{1}\}$. Similarly, she cannot pick $f_{1,L}$. If she picks some vertex in $V\setminus\{t_{1,R},f_{1,R},t_{1,L},f_{1,L}\}$, then Left’s answer is not forced, so he can pick $f_{1,R}$ to win with the same double threat as before (unless Right has picked $v_{1}$ specifically, in which case he picks $t_{1,R}$ instead and also has a double threat with $u_{1}$). This shows that Right is forced to pick $t_{1,R}$ or $f_{1,R}$ as her first move. By symmetry, assume Right has picked $t_{1,R}$ (Figure 8 with $i=1$ actually illustrates both cases). A sequence of four forced moves is now triggered:
    • Left is forced to pick $f_{1,L}$ because of the red edge $\{t_{1,R},f_{1,L}\}$.
    • Right is forced to pick $v_{1}$ because of the blue edge $\{f_{1,L},v_{1}\}$.
    • Left is forced to pick $t_{1,L}$ because of the red edge $\{t_{1,R},t_{1,L},v_{1}\}$.
    • Right is forced to pick $u_{1}$ because of the blue edge $\{t_{1,L},u_{1}\}$.
    Figure 8: For odd $i$, the updated edges induced by $V_{i}$ after the first $5(i-1)$ moves. The red (resp. blue) vertices are those that are thereafter picked by Right (resp. Left), when Right chooses to pick $t_{i,R}$ (left) or $f_{i,R}$ (right). For clarity, edges of size 2 are represented by simple lines.
    After these first five moves, all red edges and blue edges induced by $V_{1}$ are dead. Moreover, each edge of the form $e^{*}$ in the definitions of $E_{R}^{(2)}$ and $E_{L}^{(2)}$ has now yielded the updated edge $e$ since Right has picked one of $t_{1,R}$ or $f_{1,R}$ and Left has picked one of $t_{1,L}$ or $f_{1,L}$ (in fact, he has picked both).
  2. 2)
[…]

### v2

@@ 4.4 AchievementPos(3,3) is PSPACE-complete
Input : A logic formula $\phi$ in CNF form, with clauses of size exactly $3$, and an even number of variables with a prescribed ordering $x_{1},y_{1},\ldots,x_{n},y_{n}$.
Output : T if Satisfier has a winning strategy as the second player, F otherwise.
Let $\phi$ be a logic formula in CNF form, with $m$ clauses $c_{1},\ldots,c_{m}$ of size exactly $3$, and $2n$ variables with a prescribed ordering $x_{1},y_{1},\ldots,x_{n},y_{n}$. For all $j\in\llbracket 1,m\rrbracket$, write $c_{j}=\ell_{j}^{1}\vee\ell_{j}^{2}\vee\ell_{j}^{3}$, where the literals are ordered the same as their corresponding variables. We are now going to define an achievement positional game $\mathcal{G}$, with edges of size $2$ or $3$ and the same edges of size $3$ for both colors, on which Left has a winning strategy as the first player if and only if Falsifier has a winning strategy for 3-QBF on $\phi$ as the first player.
4.4.1 Idea of the construction
We simulate the 3-QBF game, with Left acting as Falsifier and Right acting as Satisfier. The game unfolds in two parts. The first part corresponds to the players choosing truth values for the variables. The second part will see Left win if and only if the valuation built during the first part satisfies $\phi$.
The first part features $2n$ sequences of moves, where each sequence of moves in $\mathcal{G}$ corresponds to a single move in the 3-QBF game. These sequences alternate between two types, depending on whether the player having to make a choice is Left (corresponding to Falsifier choosing a truth value for some $x_{i}$) or Right (corresponding to Satisfier choosing a truth value for some $y_{i}$).
Consider a sequence where Left has a choice to make. Left plays the first move, for which he only has two viable options, corresponding to his choice of a truth value for $x_{i}$. Past that initial move, guide-edges will force both players’ every move until the whole variable gadget associated with $x_{i}$ is cleared. Left plays the last move of that sequence, so that Right will play first in the next variable gadget (associated with $y_{i}$).
Sequences where Right has a choice to make work in analogous fashion, except that we need to add additional trap-edges to prevent Right from playing other moves than the prescribed ones. Some of these trap-edges have been “activated” by Left at the end of the previous sequence, producing a threat of a two-move win for Left. Therefore, Right is forced to make immediate threats herself (which she can only achieve by picking one of the expected possible choices) until she deactivates these trap-edges, which she does with the very last move of that sequence.
At the end of those $2n$ sequences of moves, a full valuation of the variables has been built, and the second part of the game starts with Left next to play. His move is actually forced by a couple of switch-edges, so that Right is next to play. Right then destroys all the clause gadgets corresponding to clauses in which at least one literal is true. For this, she plays greedy moves (as per Lemma 3.8) in the link-edges, which connect each clause gadget to its three associated variable gadgets, and then in the destruction-edges inside the clause gadgets. After she is done, Right surrenders the initiative to Left due to a couple more switch-edges. If the valuation built by the players satisfies $\phi$, then Right has destroyed all the clause gadgets, so she draws the game. Otherwise, at least one clause gadget has surviving clause-edges, which Left uses to win the game.
4.4.2 Definition of the game
Notation 4.9.
If $X$ and $Y$ are sets of sets, we define $X\otimes Y=\left\{e_{X}\cup e_{Y}\mid e_{X}\in X,e_{Y}\in Y\right\}$. We will use this notation to make edge sets easier to read in the construction, as in $\left\{\left\{v\right\}\right\}\otimes E$ where $v$ is a vertex and $E$ is an edge set can be seen as “extending every edge from $E$ with the additional vertex vv”. To alleviate notations, a set $\left\{\left\{v\right\}\right\}\otimes E$ will simply be written as $v\otimes E$.
The game $\mathcal{G}=(V,E_{L},E_{R})$ is defined as follows:
  • $V=\bigcup_{1\leq i\leq n}(V_{x_{i}}\cup V_{y_{i}})\cup\bigcup_{1\leq j\leq m}C_{j}\cup\Omega$;
  • $E_{L}=E_{1}\cup E_{LR}$;
  • $E_{R}=F_{1}\cup E_{LR}$;
where $E_{1}$ is the set of blue 2-edges, $F_{1}$ is the set of red 2-edges, and $E_{LR}$ is the set of 3-edges (all of which are both blue and red).
For vertices, we define the following sets.
  • For all $i\in\llbracket 1,n\rrbracket$, the vertices from the variable gadgets associated with the variables $x_{i}$ and $y_{i}$ respectively:
    $V_{x_{i}}=\left\{a_{i},a^{\prime}_{i},b_{i},b^{\prime}_{i},c_{i},c^{\prime}_{i},c^{\prime\prime}_{i},d_{i},f_{i},g_{i}\right\},$
    $V_{y_{i}}=\left\{r_{i},r^{\prime}_{i},s_{i},s^{\prime}_{i},t_{i},t^{\prime}_{i},t^{\prime\prime}_{i},u_{i},v_{i},w_{i}\right\}.$
    We also associate each literal $\ell$ with an index $\textup{ind}(\ell)$ defined as the unique $i\in\llbracket 1,n\rrbracket$ such that $\ell\in\left\{x_{i},\neg x_{i},y_{i},\neg y_{i}\right\}$, as well as a vertex $v(\ell)$ defined by: $v(x_{i})=b_{i}$, $v(\neg x_{i})=b^{\prime}_{i}$, $v(y_{i})=r_{i}$, $v(\neg y_{i})=r^{\prime}_{i}$, for all $i\in\llbracket 1,n\rrbracket$.
  • For all $j\in\llbracket 1,m\rrbracket$, the vertices from the clause gadget associated with the clause $c_{j}$:
    $C_{j}=\left\{\alpha_{j,1},\alpha_{j,2},\alpha_{j,3},\beta_{j,1},\beta_{j,2},\beta_{j,3}\right\}.$
  • Some additional vertices that will be important in the endgame:
    $\Omega=\left\{\omega,\omega^{\prime},\omega^{\prime\prime},\zeta,\zeta^{\prime},\zeta^{\prime\prime}\right\}.$
For edges, we define:
$E_{LR}=\bigcup_{1\leq i\leq n}(G_{i}\cup P_{i}\cup Q_{i}\cup R_{i})\cup\bigcup_{2\leq i\leq n}(E_{i}\cup F_{i})\cup\bigcup_{1\leq j\leq m}\left(L_{j}\cup D_{j}\cup\bigcup_{1\leq i\leq n}T_{i,j}\cup\Gamma_{j}\right)\cup S,$
and the following sets comprising six different types of edges.
  • The guide-edges, which force moves in the variable gadgets, starting with the guide-edges for Left’s first choice:
    • $E_{1}=\left\{\left\{a_{1},b_{1}\right\},\left\{a^{\prime}_{1},b^{\prime}_{1}\right\}\right\}$;
    • $F_{1}=\left\{\left\{b_{1},c_{1}\right\},\left\{b^{\prime}_{1},c_{1}\right\},\left\{c_{1},c^{\prime}_{1}\right\},\left\{c_{1},c^{\prime\prime}_{1}\right\},\left\{d_{1},f_{1}\right\},\left\{a_{1},f_{1}\right\},\left\{a^{\prime}_{1},f_{1}\right\}\right\}$;
    • $G_{1}=\left\{\left\{a_{1},c_{1},d_{1}\right\},\left\{a^{\prime}_{1},c_{1},d_{1}\right\},\left\{a_{1},d_{1},g_{1}\right\},\left\{a^{\prime}_{1},d_{1},g_{1}\right\}\right\}$;
    for all $i\in\llbracket 1,n\rrbracket$, the guide-edges for Right’s $i$-th choice:
    • $P_{i}=f_{i}\otimes\left\{\left\{s_{i},t_{i}\right\},\left\{s^{\prime}_{i},t_{i}\right\},\left\{t_{i},t^{\prime}_{i}\right\},\left\{t_{i},t^{\prime\prime}_{i}\right\},\left\{u_{i},v_{i}\right\},\left\{r_{i},v_{i}\right\},\left\{r^{\prime}_{i},v_{i}\right\}\right\};$
    • $Q_{i}=d_{i}\otimes\left\{\left\{r_{i},s_{i}\right\},\left\{r^{\prime}_{i},s^{\prime}_{i}\right\}\right\};$
    • $R_{i}=\left\{\left\{r_{i},t_{i},u_{i}\right\},\left\{r^{\prime}_{i},t_{i},u_{i}\right\},\left\{s_{i},s^{\prime}_{i},w_{i}\right\}\right\};$
    and for all $i\in\llbracket 2,n\rrbracket$, the guide-edges for Left’s $i$-th choice:
    • $E_{i}=u_{i-1}\otimes\left\{\left\{a_{i},b_{i}\right\},\left\{a^{\prime}_{i},b^{\prime}_{i}\right\}\right\};$
    • $F_{i}=v_{i-1}\otimes\left\{\left\{b_{i},c_{i}\right\},\left\{b^{\prime}_{i},c_{i}\right\},\left\{c_{i},c^{\prime}_{i}\right\},\left\{c_{i},c^{\prime\prime}_{i}\right\},\left\{d_{i},f_{i}\right\},\left\{a_{i},f_{i}\right\},\left\{a^{\prime}_{i},f_{i}\right\}\right\};$
    • $G_{i}=\left\{\left\{a_{i},c_{i},d_{i}\right\},\left\{a^{\prime}_{i},c_{i},d_{i}\right\},\left\{a_{i},d_{i},g_{i}\right\},\left\{a^{\prime}_{i},d_{i},g_{i}\right\}\right\}.$
  • For all $j\in\llbracket 1,m\rrbracket$, the link-edges, which connect the clause gadgets to their associated variable gadgets:
    $L_{j}=\left\{\{v(\ell_{j}^{k}),\alpha_{j,k},\beta_{j,k}\}\mid k\in\llbracket 1,3\rrbracket\right\}.$
  • For all $j\in\llbracket 1,m\rrbracket$, the destruction-edges, which ensure that a clause gadget can be completely destroyed by Right as long as at least one literal from the corresponding clause is true:
    $D_{j}=\left\{\left\{\alpha_{j,1},\beta_{j,1},\alpha_{j,3}\right\},\left\{\alpha_{j,2},\beta_{j,2},\alpha_{j,1}\right\},\left\{\alpha_{j,3},\beta_{j,3},\alpha_{j,2}\right\}\right\},$
  • For all $(i,j)\in\llbracket 1,n\rrbracket\times\llbracket 1,m\rrbracket$, the trap-edges, which prevent Right from playing some moves too early:
    $T_{i,j}=\left\{\left\{f_{i},v_{i}\right\}\right\}\otimes\left\{\left\{\alpha_{j,1}\right\},\left\{\beta_{j,1}\right\},\left\{\alpha_{j,2}\right\},\left\{\beta_{j,2}\right\},\left\{\alpha_{j,3}\right\},\left\{\beta_{j,3}\right\}\right\}.$
  • For all $j\in\llbracket 1,m\rrbracket$, the clause-edges associated with the clause $c_{j}$:
    $\Gamma_{j}=\,\omega\otimes\left\{\left\{\alpha_{j,1},\alpha_{j,2}\right\},\left\{\alpha_{j,2},\alpha_{j,3}\right\},\left\{\alpha_{j,3},\alpha_{j,1}\right\}\right\}.$
  • The switch-edges, which allow us to control which player is next to play at two key moments during the endgame, by forcing the opponent to waste a move:
    $S=\,v_{n}\otimes\omega\otimes\left\{\left\{\omega^{\prime}\right\},\left\{\omega^{\prime\prime}\right\}\right\}\,\cup\,\omega\otimes\zeta\otimes\left\{\left\{\zeta^{\prime}\right\},\left\{\zeta^{\prime\prime}\right\}\right\}.$
In all subsequent figures, every (updated) red (resp. blue) 2-edge will be represented as a red (resp. blue) line joining its two vertices. Although all 3-edges are both blue and red in the game, we represent them using various colors to help readability.
The variable gadgets are represented in Figures 9 and 10. The clause gadgets are represented in Figure 11.
Figure 9: Gadget associated with the variable $x_{i}$, as updated just before play starts inside it, i.e. at the end of Phase i1i-1 of regular play. For $i=1$, the gadget is like this at the start of the game.
Figure 10: Gadget associated with the variable $y_{i}$, as updated just before play starts inside it, i.e. after Steps (1) and (2) of Phase $i$ of regular play.
Figure 11: Clause gadget associated with the clause $c_{j}=\ell_{j}^{1}\vee\ell_{j}^{2}\vee\ell_{j}^{3}$. The vertices $f_{i}$ and $v_{i}$ are represented multiple times for readability. The three link-edges are in green. The three destruction-edges are in purple. The three clause-edges are in yellow. The $6n$ trap-edges are in orange.
4.4.3 Regular play
Assuming Left starts, we define regular play on $\mathcal{G}$, during which the players build a valuation $\mu$ for the variables $x_{1},y_{1}\ldots,x_{n},y_{n}$. An arrow labelled “f” signifies a forced move (when a player cannot win in one move but the opponent threatens to do so), while an arrow labelled “g” signifies a greedy move in the sense of Lemma 3.8. For $i$ from $1$ to $n$, Phase $i$ of regular play consists of the following four steps. Steps (1) and (2) take place inside the variable gadget associated with $x_{i}$, with the situation at the start of Step (1) as shown in Figure 9. Steps (3) and (4) take place inside the variable gadget associated with $y_{i}$, with the situation at the start of Step (3) as shown in Figure 10.
  1. (1)
    Left chooses the value T or F for $x_{i}$, by picking $a_{i}$ or $a^{\prime}_{i}$ respectively. Both options trigger a sequence of forced moves inside the associated variable gadget:
[…]
  2. (2)
    Right now has no forced move, and uses this window to play a greedy move. Either $a_{i}$ or $a^{\prime}_{i}$ has not been played, and is a greedy move for Right.
    $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a^{\prime}_{i}$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}g_{i}$}}gf
    or   $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{i}$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}g_{i}$}}gf
  3. (3)
    Right chooses the value T or F for $y_{i}$, by picking $r_{i}$ or $r^{\prime}_{i}$ respectively. Both options trigger a sequence of forced moves inside the associated variable gadget:
[…]
  4. (4)
    Left now has no forced move, and uses this window to play a greedy move. Either $s_{i}$ or $s^{\prime}_{i}$ has not been played, and is a greedy move for Left.
    $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}s^{\prime}_{i}$$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w_{i}$}}gf
    or   $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}s_{i}$$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w_{i}$}}gf
We now list some important properties of regular play. We say a blue (resp. red) edge is dead if Right (resp. Left) has picked at least one of its vertices. Dead edges are those that have yielded no edge in the updated game. We say an edge is intact if none of its vertices has been picked.
Claim 1.
If both players obey regular play then, for all $i\in\llbracket 0,n\rrbracket$, we have the following at the end of Phase $i$ (end of Phase 0 being the beginning of the game):
  1. (i)
    All picked vertices are in $\bigcup_{1\leq t\leq i}(V_{x_{t}}\cup V_{y_{t}})$. In particular, all destruction-edges and clause-edges are intact.
  2. (ii)
    For all $t\in\llbracket 1,i\rrbracket$, every guide-edge in $E_{t}\cup F_{t}\cup G_{t}\cup P_{t}\cup Q_{t}\cup R_{t}$ is dead.
  3. (iii)
    For all $t\in\llbracket 1,i\rrbracket$, Left has picked $f_{t}$ and $u_{t}$ while Right has picked $d_{t}$ and $v_{t}$, so in particular every trap-edge in $\bigcup_{1\leq j\leq m}T_{t,j}$ is dead.
  4. (iv)
    For all $(j,k)\in\llbracket 1,m\rrbracket\times\llbracket 1,3\rrbracket$, the vertex $v(\ell_{j}^{k})$ has been picked by Right if $\textup{ind}(\ell_{j}^{k})\leq i$ and $\mu(\ell_{j}^{k})={\sf T}$ or has not been picked at all otherwise, so in particular the link-edge $\{v(\ell_{j}^{k}),\alpha_{j,k},\beta_{j,k}\}$ has yielded the updated red edge $\left\{\alpha_{j,k},\beta_{j,k}\right\}$ if $\textup{ind}(\ell_{j}^{k})\leq i$ and $\mu(\ell_{j}^{k})={\sf T}$ or is intact otherwise. Here, $\mu$ refers to the partial valuation that the players have built so far (valuation of $x_{1},y_{1},\ldots,x_{i},y_{i}$).
  5. (v)
    The switch-edges are intact, except if $i=n$ for the two switch-edges $v_{n}\otimes\omega\otimes\left\{\left\{\omega^{\prime}\right\},\left\{\omega^{\prime\prime}\right\}\right\}$ which have yielded the updated red edges $\left\{\omega,\omega^{\prime}\right\}$ and $\left\{\omega,\omega^{\prime\prime}\right\}$.
Proof of Claim 1.
We prove the claim by induction on $i$. The case $i=0$ is trivially true. Now, let $i\in\llbracket 1,n\rrbracket$ such that items (i)–(v) hold for i1i-1. Recall that, by definition of regular play, during Phase $i$:
  • Either Left chooses $\mu(x_{i})={\sf T}$, in which case he picks $a_{i},c_{i},f_{i},g_{i}$ and Right picks $b_{i},d_{i},a^{\prime}_{i}$, or Left chooses $\mu(x_{i})={\sf F}$, in which case he picks $a^{\prime}_{i},c_{i},f_{i},g_{i}$ and Right picks $b^{\prime}_{i},d_{i},a_{i}$.
  • Then, either Right chooses $\mu(y_{i})={\sf T}$, in which case she picks $r_{i},t_{i},v_{i},w_{i}$ and Left picks $s_{i},u_{i},s^{\prime}_{i}$, or Right chooses $\mu(x_{i})={\sf F}$, in which case she picks $r^{\prime}_{i},t_{i},v_{i},w_{i}$ and Left picks $s^{\prime}_{i},u_{i},s_{i}$.
We now show that items (i)–(v) hold after these moves have been played.
  1. (i)
    By the induction hypothesis, every vertex that was picked during Phases $1$ to i1i-1 is in $\bigcup_{1\leq t\leq i-1}(V_{x_{t}}\cup V_{y_{t}})$. Moreover, every vertex that was picked during Phase $i$ is in $V_{x_{i}}\cup V_{y_{i}}$. The consequence on the destruction-edges and clause-edges is clear, since none of them intersects $\bigcup_{1\leq t\leq n}(V_{x_{t}}\cup V_{y_{t}})$.
  2. (ii)
    Using the induction hypothesis, we only need to address the case $t=i$. By symmetry, assume that the chosen values are $\mu(x_{i})={\sf T}$ and $\mu(y_{i})={\sf T}$. This means that $a_{i},c_{i},f_{i},g_{i},s_{i},u_{i},s^{\prime}_{i}$ have been picked by Left, and $b_{i},d_{i},a^{\prime}_{i},r_{i},t_{i},v_{i},w_{i}$ have been picked by Right. Item (iii) of the induction hypothesis ensures that, before the start of Phase $i$, Left had picked $u_{i-1}$ and that Right had picked $v_{i-1}$. Knowing this, we can easily check that every edge in $E_{i}\cup F_{i}\cup G_{i}\cup P_{i}\cup Q_{i}\cup R_{i}$ is dead:
    • Every edge in $E_{i}$ contains $u_{i-1}$ (picked by Left). They also all contain one of $b_{i}$ or $a^{\prime}_{i}$ (both picked by Right). All in all, every edge in $E_{i}$ is dead by the end of Phase $i$. The case of $Q_{i}$ is analogous.
    • Every edge in $F_{i}$ contains one of $c_{i}$ or $f_{i}$ (both picked by Left), and also contains $v_{i-1}$ (picked by Right), so it is dead by the end of Phase $i$. The case of $P_{i}$ is analogous.
    • Every edge in $G_{i}$ contains one of $a_{i}$, $c_{i}$ or $g_{i}$ (all picked by Left), and one of $d_{i}$ or $a^{\prime}_{i}$ (both picked by Right), so it is dead by the end of Phase $i$.
    • Every edge in $R_{i}$ contains one of $u_{i}$ or $s_{i}$ (both picked by Left), and one of $t_{i}$ or $w_{i}$ (both picked by Right), so it is dead by the end of Phase $i$.
  3. (iii)
    Using the induction hypothesis, we only need to address the case $t=i$. Clearly, Left has picked $f_{i}$ and $u_{i}$, while Right has picked $d_{i}$ and $v_{i}$. Every trap-edge in $T_{i,j}$ contains both $f_{i}$ and $v_{i}$, and is therefore dead.
  4. (iv)
    Using the induction hypothesis, we only need to address the case $\textup{ind}(\ell_{j}^{k})=i$, i.e. $\ell_{j}^{k}\in\left\{x_{i},\neg x_{i},y_{i},\neg y_{i}\right\}$.
    • Suppose $\ell_{j}^{k}=x_{i}$, i.e. $v(\ell_{j}^{k})=b_{i}$. If $\mu(\ell_{j}^{k})={\sf T}$, then $\mu(x_{i})={\sf T}$, so by definition of regular play $b_{i}$ has been picked by Right. If $\mu(\ell_{j}^{k})={\sf F}$, then $\mu(x_{i})={\sf F}$, so by definition of regular play $b_{i}$ has not been picked.
    • Suppose $\ell_{j}^{k}=\neg x_{i}$, i.e. $v(\ell_{j}^{k})=b^{\prime}_{i}$. If $\mu(\ell_{j}^{k})={\sf T}$, then $\mu(x_{i})={\sf F}$, so $b^{\prime}_{i}$ has been picked by Right. If $\mu(\ell_{j}^{k})={\sf F}$, then $\mu(x_{i})={\sf T}$, so $b^{\prime}_{i}$ has not been picked.
    • Suppose $\ell_{j}^{k}=y_{i}$, i.e. $v(\ell_{j}^{k})=r_{i}$. If $\mu(\ell_{j}^{k})={\sf T}$, then $\mu(y_{i})={\sf T}$, so by definition of regular play $r_{i}$ has been picked by Right. If $\mu(\ell_{j}^{k})={\sf F}$, then $\mu(y_{i})={\sf F}$, so by definition of regular play $b_{i}$ has not been picked.
    • Finally, suppose $\ell_{j}^{k}=\neg y_{i}$, i.e. $v(\ell_{j}^{k})=r^{\prime}_{i}$. If $\mu(\ell_{j}^{k})={\sf T}$, then $\mu(y_{i})={\sf F}$, so $r^{\prime}_{i}$ has been picked by Right. If $\mu(\ell_{j}^{k})={\sf F}$, then $\mu(y_{i})={\sf T}$, so $r^{\prime}_{i}$ has not been picked.
    The consequence on the link-edges is clear.
  5. (v)
    This is a consequence of items (i) and (iii). If $i=n$, then Right has picked $v_{n}$, hence the updated red edges $\left\{\omega,\omega^{\prime}\right\}$ and $\left\{\omega,\omega^{\prime\prime}\right\}$.∎
A consequence of Claim 1 is that regular play ends without a winner. We now explain what happens afterwards.
Claim 2.
Assume both players are constrained to obey regular play until it ends. Then, Left has a winning strategy on $\mathcal{G}$ as the first player if and only if Falsifier has a winning strategy for 3-QBF on $\phi$ as the first player.
Proof of Claim 2.
Consider the situation at the end of regular play, as summarized by Claim 1 for $i=n$. All guide-edges and trap-edges are dead. Some link-edges and switch-edges have yielded updated red 2-edges. All other edges are intact. In particular, all updated blue edges have size 3. Therefore, Left is now forced to destroy the red $P_{3}$ formed by the updated red edges $\{\omega,\omega^{\prime}\}$ and $\{\omega,\omega^{\prime\prime}\}$, otherwise Right would pick $\omega$ next and win one move later by picking $\omega^{\prime}$ or $\omega^{\prime\prime}$ (see Figure 12, left). Since these two are the only updated edges containing $\omega^{\prime}$ and $\omega^{\prime\prime}$ respectively, Lemma 3.6 ensures that it is optimal for Left to pick $\omega$. We assume that Left does pick $\omega$.
Figure 12: The switch-edges, as updated at the end of regular play (on the left) and after Left picks $\omega$ (on the right).
Right is now next to play, and her move is not forced. Indeed, even though the two switch-edges $\omega\otimes\zeta\otimes\left\{\left\{\zeta^{\prime}\right\},\left\{\zeta^{\prime\prime}\right\}\right\}$ have yielded a blue $P_{3}$ in the updated game due to Left picking $\omega$ (see Figure 12, right), Right does have some updated red 2-edges of her own.
[…]